Diffusion Morning Report 05/31/05

Ted Berg tberg at snf.stanford.edu
Tue May 31 06:48:22 PDT 2005


Good Morning All,
    Looking  ok this morning.  Epi has facilities problems-checking it 
out. Nitride is having intermittant gate valve issues-looking into it.
Tylan 1 is still down. LTO is down for pull and clean. Everything else 
looks ok for now.

  	Date: 	5/26/2005 	THICKNESS-Change at
Tube Name/Number 	Status 	Comments 	 
Tylan 1 	Down 	Comm failure 	 
Tylan 2 	Up 	  	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-TEOS2 	Up 	  	 
Tylan 8-SiGe 	Up 	  	3.12/30
Tylan 9-Poly 	Up 	  	1.84/30
Tylan 10-Nitride 	yellow 	Particles look good 	13.39/30
Tylan 11-BPSG-LTO 	Down 	pull and clean 	29.12/30
Tylan 12-FGA 	Up 	  	 
Tystar 	Up 	  	3.58/30
Doped Poly 	GONE 	  	 
TEOS 	yellow 	  	 
Tylan 16 	N/A 	  	 
Thermco 1 	Up 	  	 
Thermco 2 	Up 	  	 
Thermco Poly 	yellow 	testing doped 	 
Thermco Nitride 	Down 	Waiting 	 
  	  	  	 
Burn Box 	Up 	  	 
ZYGO 	up 	  	.
  	  	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
  	  	  	 
SiH4 for furnaces 	  	B on line 280 PSI 	 changeover working well 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20050531/f29c92d1/attachment.html>


More information about the diffusion mailing list